# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ 2.4 Additional System Variables and Operational Constraints
The mathematical modeling of the operation of the P2H system, as shown in Figures 1 and 2, requires the consideration of several additional variables that have not been included in the set of state and control variables. They are referred to as system variables and are subject to certain operational constraints, which are explained in this subsection. These system variables and their dynamics are needed to derive state and state-dependent control constraints of the control problem that we formulate in Section 3. However, the system variables are “internal” variables whose specific values do not need to be observed by the controller and which are not included in the decision-making process. The latter is based solely on knowledge of the state variables.
To formulate the mathematical model discussed in this work, we make the following simplifying assumption
@@
The mass flow $\dot{m}$ of the thermal oil stream and the temperature $T^{\text{LT,in}}$ of the waste heat air stream at the LTHX inlet are constant over the entire period $[0,t_{\text{E}}]$.
@@ 2.4.1 Steam Generator
The nonlinear functions $F^{\text{SG,in}}$ and $F^{\text{SG,out}}$ represent surrogate models for the underlying energy balances and are generated using process simulations, see Appendix B. According to Assumption 2.2, the mass flow $\dot{m}$ is constant in our model, so $T^{\text{SG,in}}$ and $T^{\text{SG,out}}$ are also constant over the entire period $[0,t_{\text{E}}]$. It is obvious that $T^{\text{SG,in}}>T^{\text{SG,out}}$, because the SG can simply be understood as a heat exchanger to supply the factory with superheated steam.
@@ 2.4.2 High-Temperature Heat Pump
According to Assumption 2.1, the control variable $A$ representing the heat flow rate in the TES is piecewise constant, i.e. $A(t)=A_{n}$ in each interval $[t_{n},t_{n+1})$ for $n=0,\ldots,N-1$. This property is transferred to the HTHX inlet and outlet temperatures $T^{\text{HT,in}}$ and $T^{\text{HT,out}}$, which results from the given system configuration in which the TES is integrated, the fact that $T^{\text{SG,in}}$ and $T^{\text{SG,out}}$ are constant on $[0,t_{\text{E}}]$ and the following energy balances.
In period $[t_{n},t_{n+1})$ during charging, we have an inflow of thermal energy from the HTHP to the thermal oil loop with rate $n_{\text{H}}\dot{m}c_{F}T^{\text{HT,out}}(t)$ and an outflow from this loop to the SG with constant rate $n_{\text{H}}\dot{m}c_{F}T^{\text{SG,in}}$. Here, $c_{F}$ denotes the specific heat capacity (assumed to be temperature independent) of the thermal oil and $n_{\text{H}}$ the number of HTHPs operating in parallel in the underlying P2H system, which is not explicitly shown in Figure 2). Since we neglect losses to the environment, the difference between the two rates gives the constant inflow rate $A_{n}$ to the TES during the charging process. Recall that it holds $A_{n}>0$ in charging mode, thus it follows $T^{\text{HT,out}}>T^{\text{SG,in}}$.
Analogously, during discharging, there is an inflow of thermal energy from the SG with constant rate $n_{\text{H}}\dot{m}c_{F}T^{\text{SG,out}}$ and an outflow to the HTHP with rate $n_{\text{H}}\dot{m}c_{F}T^{\text{HT,in}}(t)$. The difference of the two rates determines the outflow rate from the TES, which is $A_{n}$. Since $A_{n}<0$ during discharging, it follows $T^{\text{HT,in}}>T^{\text{SG,out}}$.
@@
$\begin{split}A_{n}&=n_{\text{H}}\dot{m}\,c_{F}(T^{\text{HT,out}}(t)-T^{\text{SG,in}})>0\quad\text{during charging, ~{}and }\\ A_{n}&=n_{\text{H}}\dot{m}\,c_{F}(T^{\text{SG,out}}-T^{\text{HT,in}}(t))<0\quad\text{during discharging,}\end{split}$ (5)
@@
$\begin{array}[]{rllll}T^{\text{HT,out}}(t)&=T^{\text{HT,out}}_{n}&=\tau^{\text{out}}(A_{n})&\text{with}\quad\tau^{\text{out}}(a)&=T^{\text{SG,in}}+\frac{a^{+}}{n_{\text{H}}\dot{m}\,c_{F}},\quad\text{ and}\\[4.30554pt] T^{\text{HT,in}}(t)&=T^{\text{HT,in}}_{n}&=\tau^{\text{in}}(A_{n})&\text{with}\quad\tau^{\text{in}}(a)&=T^{\text{SG,out}}+\frac{a^{-}}{n_{\text{H}}\dot{m}\,c_{F}},\end{array}$ (8)
@@
with $g^{HF}(\tau^{\text{in}},\tau^{\text{out}})=n_{\text{H}}\dot{m}\,c_{F}(\tau^{\text{out}}-\tau^{\text{in}}-(T^{\text{SG,in}}-T^{\text{SG,out}}))$. Note that $T^{\text{HT,in}}_{n}=T^{\text{SG,out}}$ during charging, $T^{\text{HT,out}}_{n}=T^{\text{SG,in}}$ during discharging, and both equations hold true during idle periods.
@@
From Walden2023 it is known that there is a complex relationship between the HTHX oil outlet temperature $T^{\text{HT,out}}$ and the HTHX oil inlet temperature $T^{\text{HT,in}}$, the HTF mass flow $\dot{m}$, the waste heat air temperature $T^{\text{LT,in}}$ at the LTHX inlet, and the compressor shaft speed $D$, which can be expressed by a surrogate model in terms of a multivariate cubic polynomial $F_{1}$ as $T^{\text{HT,out}}(t)=F_{1}(T^{\text{HT,in}}(t),\dot{m},T^{\text{LT,in}},D(t))$. As we know from (8) that $T^{\text{HT,in}}$ and $T^{\text{HT,out}}$ are stepwise functions, i.e. $T^{\text{HT,out}}(t)=T^{\text{HT,out}}_{n}$ and $T^{\text{HT,in}}(t)=T^{\text{HT,in}}_{n}$ in each interval $[t_{n},t_{n+1})$ for $n=0,\ldots,N-1$, and based on the Assumption 2.2, it follows directly that $D$ is also constant between two subsequent discrete time points, so we obtain
@@
More specifically, the compressor shaft speed $D$ can be varied (only at discrete time grid points) to determine the pressure ratio within the HTHP and thus the thermal oil temperature $T^{\text{HT,out}}$ at the HTHX outlet. Details on the definition of the polynomial function $F_{1}$ are provided in Appendix B.
@@
Another complex relationship from Walden2023 holds for the electricity consumption $P^{\text{H}}$ of the HTHP. Again this is expressed by a surrogate model in the form of a multivariate quadratic polynomial $F_{2}$ for $n=0,\ldots,N$ as
@@
The factor $n_{\text{H}}$ in (11) as already mentioned above, denoting the number of HTHPs operating in parallel, takes into account that the surrogate model represents only a single HTHP. The detailed definition of $F_{2}$ can be found in Appendix B.
@@
$P^{\text{H}}_{n}=P^{\text{G}}(t)+P^{\text{W}}(t),\qquad t\in[t_{n},t_{n+1}),~{}n=0,\ldots,N-1.$ (12)
For $P^{\text{G}}>0$, electricity is drawn (purchased) from the grid, for $P^{\text{G}}<0$, electricity is fed (sold) into the grid. Note that the dependence of the WT power $P^{\text{W}}$ on the wind speed $W$ is described by the so-called power curve, $P^{\text{W}}=P_{\text{WT}}(W)$. This is a nonlinear function that grows cubically WANG2019109422 at medium wind speeds until it reaches the rated power. This value is kept constant for higher speeds and is set to zero above the cut-out speed and below a cut-in speed of the turbine. More details are given in Appendix E.
@@
Based on (10) and (11), the HTHP’s electricity consumption $P^{\text{H}}_{n}$ can be determined at each time point $t_{n}$ for a given control $A$, which determines by (8) the temperatures $T^{\text{HT,out}}_{n}$ and $T^{\text{HT,in}}_{n}$ of the HTF at the outlet and inlet of the HTHX, respectively. Suppose that at time $t_{n}$ the control is set to be $A_{n}=a$, (8) implies $T^{\text{HT,out}}_{n}=\tau^{\text{out}}(a)$ and $T^{\text{HT,in}}_{n}=\tau^{\text{in}}(a)$, and recall that $\dot{m}$ and $T^{\text{LT,in}}$ are constants. Then, in a first step, the corresponding shaft speed $D_{n}=d$ is determined by solving (10) for the unknown $d$, i.e. $\tau^{\text{out}}(a)=F_{1}(d):=F_{1}(\tau^{\text{in}}(a),\dot{m},T^{\text{LT,in}},d)$. Since $F_{1}$ is a cubic polynomial in $d$, the solution is among the real-valued roots of this polynomial. For the given $F_{1}$ and using the fact that the shaft speed is restricted to values within the interval $[d_{\text{min}},d_{\text{max}}]$, defined by technical conditions Walden2023 , there is a unique root $d^{*}=d^{*}(a)$ in this interval. In a second step, the corresponding electricity consumption $P^{\text{H}}_{n}$ is obtained by substituting $d^{*}$ into (11) via
@@
To emphasize the dependence of $P^{\text{H}}$ on the control $A$, we introduce the function $\pi^{\text{H}}$ that maps $A$ to the electricity consumption, i.e.
$P^{\text{H}}_{n}=\pi^{\text{H}}(A_{n}),\quad n=0,\ldots,N-1,~{}~{}\text{with}~{}\pi^{\text{H}}(a)=n_{\text{H}}F_{2}(\tau^{\text{in}}(a),\dot{m},T^{\text{LT,in}},d^{*}(a)).$ (14)
@@ 2.4.3 Thermal Energy Storage Operational Modes
$\begin{array}[]{llll}\text{Charging:}&\quad A_{n}>0,&\quad T^{\text{HT,in}}=T^{\text{SG,out}},&\quad l^{\text{C}}\in(0,1],\\ \text{Discharging:}&\quad A_{n}<0,&\quad T^{\text{HT,out}}_{n}=T^{\text{SG,in}},&\quad l^{\text{D}}\in(0,1],\\ \text{Idle:}&\quad A_{n}=0,&\quad l^{\text{C}}=l^{\text{D}}=0.&\end{array}$ (15)
In charging mode, the excess thermal energy is stored in the sensible TES and increases its medium temperature, while during discharging the cooler HTF with temperature $T^{\text{SG,out}}$ absorbs heat from the TES, and decreases the temperature of the storage medium. The heat emission and absorption of the HTF is described by its respective charging and discharging efficiencies $\varepsilon^{C},\varepsilon^{D}\in(0,1]$, which directly influence the TES outlet temperatures. For details we refer to Appendix D. In idle mode, the HTF bypasses the TES completely and obviously implies $T^{\text{HT,out}}_{n}=T^{\text{SG,in}}$ and $T^{\text{HT,in}}_{n}=T^{\text{SG,out}}$, which also follows directly from (85) and (88) in Appendix D. There, we provide also further details, including the dependence of the time-varying charging and discharging factors $l^{\text{C}},l^{\text{D}}$ on the TES temperature $R$ and the chosen control $A_{n}$.
@@ 3.1 State Dynamics
Throughout this paper all stochastic processes and random variables are supposed to be defined on a filtered probability space $(\Omega,\pazocal{F},\mathbb{F},\mathbb{P})$. In particular, that space carries a two-dimensional Brownian motion $B=(B^{W},B^{S})$, with two independent standard Brownian motions $B^{W},B^{S}$ on $[0,t_{\text{E}}]$, which will be used below to drive stochastic differential equations (SDEs) describing the dynamics of $W$ and $S$. The filtration $\mathbb{F}$ is assumed to be generated by $B$, that is, $\mathbb{F}=\mathbb{F}^{B}=(\pazocal{F}^{B}(t))_{t\in[0,t_{\text{E}}]}$ with the $\sigma$-algebras $\pazocal{F}^{B}(t)=\sigma\{B(s),s\leq t\}$, augmented by the $\mathbb{P}$-nullsets, so that $\mathbb{F}$ satisfies the usual assumptions. While, the market price of electricity $S$ can also take negative values, the wind speed $W$ is always non-negative. We therefore replace $W$ with $\log W$ and assume for $t\in[0,t_{\text{E}}]$
$\begin{split}\log W(t)&=\mu_{W}(t)+Y^{W}(t),\\ S(t)&=\mu_{S}(t)+Y^{S}(t).\end{split}$ (16)
@@
with mean reversion speeds $\lambda_{W},\lambda_{S}>0$, diffusion coefficients $\sigma_{W},\sigma_{S}>0$ and a constant $c_{W}\geq 0$. Due to the different natures of wind speed and electricity price, we assume that $\lambda_{W}\neq\lambda_{S}$ to simplify our analysis. A positive constant $c_{W}$ leads to a negative correlation of the wind speed $W$ and the price process $S$, see below in Lemma 3.1 and Proposition 3.2. This is often observed in energy markets.
The two SDEs above have analytical solutions that allow us to derive the following closed-form expressions for the joint distribution of the pair of random variables $W$ and $S$ as follows, with the proofs provided in Appendix E.1.
@@
Let $0\leq t_{a}<t_{b}\leq t_{\text{E}}$. Then the solutions of the SDEs (17) on $[t_{a},t_{b}]$ to given initial values $Y^{W}(t_{a})=y_{W}$ and $Y^{S}(t_{a})=y_{S}$ with $y_{W},y_{S}\in\mathbb{R}$ are Ornstein-Uhlenbeck processes with
@@
Combining the above result for $t_{a}=t_{n}$ and $t_{b}=t_{n+1}=t_{a}+\Delta t$ with the definitions from (16), and recall the notation $W_{n}=W(t_{n}),S_{n}=S(t_{n})$ for $n=0,\ldots,N$, we obtain the following result for the joint distribution of $(\log W_{n+1},S_{n+1})$ given $(\log W_{n},S_{n})$. This will be useful for the construction of the transition operator and the associated transition kernel for the MDP’s state process below in (36).
@@
$m^{W}_{n+1}(w)$ $=\mu_{W}(t_{n+1})+m_{Y^{W}}(\Delta t,\log w-\mu_{W}(t_{n})),$ (23)
$m^{S}_{n+1}(w,s)$ $=\;\mu_{S}(t_{n+1})\,+m_{Y^{S}}(\Delta t,\log w-\mu_{W}(t_{n}),s-\mu_{S}(t_{n})),$ (24)
where $m_{Y^{W}},m_{Y^{S}}$ and $\Sigma_{Y}$ are given in Lemma 17
Note that the wind speed $W$ follows a log-normal distribution because $\log W$ is Gaussian. The above result on the conditional distribution of the pairs $(\log W_{n},S_{n})$ and the fact that the dynamics of the stochastic fluctuations $Y^{W},Y^{S}$ are driven by Brownian motions, i.e. processes with independent increments, can be used to derive a recursion for the discrete-time dynamics of the sequence of these pairs which is driven by a sequence of independent standard normally distributed random vectors.

### v3

@@ 2.4 Additional System Variables and Operational Constraints
The mathematical modeling of the operation of the P2H system, as shown in Figures 1 and 2, requires the consideration of several additional variables that have not been included in the set of state and control variables. They are referred to as system variables and are subject to certain operational constraints, which are explained in this subsection. These system variables and their dynamics are needed to derive state-dependent control constraints of the control problem that we formulate in Section 3. However, the system variables are regarded as internal variables whose specific values do not need to be observed by the controller and which are not included in the decision-making process. The latter is based solely on knowledge of the state variables.
To formulate the mathematical model discussed in this work, we make the following simplifying
@@
The mass flow $\dot{m}$ of the thermal oil stream and the temperature $T^{\text{LT,in}}$ of the waste heat air stream at the LTHX inlet are constant over the entire period $0,t_{\text{E}}]$.
@@ 2.4.1 Steam Generator
The nonlinear functions $F^{\text{SG,in}}$ and $F^{\text{SG,out}}$ represent surrogate models for the underlying energy balances and are generated using process simulations, see Section Appendix C. According to Assumption 2.2, the mass flow $\dot{m}$ is constant in our model, so $T^{\text{SG,in}}$ and $T^{\text{SG,out}}$ are also constant over the entire period $0,t_{\text{E}}]$. It is obvious that $T^{\text{SG,in}}>T^{\text{SG,out}}$, because the SG can simply be understood as a heat exchanger to supply the factory with superheated steam.
@@ 2.4.2 High-Temperature Heat Pump
According to Assumption 2.1, the control variable $A$ representing the heat flow rate in the TES is piecewise constant, i.e., $A(t)=A_{n}$ in each interval $t_{n},t_{n+1})$ for $n=0,\ldots,N-1$. This property is transferred to the HTHX inlet and outlet temperatures $T^{\text{HT,in}}$ and $T^{\text{HT,out}}$, which result from the given system configuration in which the TES is integrated, the fact that $T^{\text{SG,in}}$ and $T^{\text{SG,out}}$ are constant on $0,t_{\text{E}}]$ and the following energy balances.
In period $t_{n},t_{n+1})$ during charging, we have an inflow of thermal energy from the HTHP to the thermal oil loop with rate $n_{\text{H}}\dot{m}c_{\text{p,f}}T^{\text{HT,out}}(t)$ and an outflow from this loop to the SG with constant rate $n_{\text{H}}\dot{m}c_{\text{p,f}}T^{\text{SG,in}}$. Here, $c_{\text{p,f}}$ denotes the specific heat capacity (assumed to be temperature independent) of the thermal oil and $n_{\text{H}}$ the number of HTHPs operating in parallel in the underlying P2H system, which is not explicitly shown in Figure 2). Since we neglect losses to the environment, the difference between the two rates gives the constant inflow rate $A_{n}$ to the TES during the charging process. Recall that $A_{n}>0$ holds in charging mode, thus it follows $T^{\text{HT,out}}>T^{\text{SG,in}}$.
Analogously, during discharging, there is an inflow of thermal energy from the SG with constant rate $n_{\text{H}}\dot{m}c_{\text{p,f}}T^{\text{SG,out}}$ and an outflow to the HTHP with rate $n_{\text{H}}\dot{m}c_{\text{p,f}}T^{\text{HT,in}}(t)$. The difference of the two rates determines the outflow rate from the TES, which is $A_{n}$. Since $A_{n}<0$ during discharging, it follows $T^{\text{HT,in}}>T^{\text{SG,out}}$.
@@
$\begin{split}A_{n}&=n_{\text{H}}\dot{m}\,c_{\text{p,f}}(T^{\text{HT,out}}(t)-T^{\text{SG,in}})>0,\quad\text{during charging ~and }\\ A_{n}&=n_{\text{H}}\dot{m}\,c_{\text{p,f}}(T^{\text{SG,out}}-T^{\text{HT,in}}(t))<0,\quad\text{during discharging}\end{split}$ (5)
@@
$\begin{array}[]{rllll}T^{\text{HT,out}}(t)&=T^{\text{HT,out}}_{n}&=\tau^{\text{out}}(A_{n})&\text{with}\quad\tau^{\text{out}}(a)&=T^{\text{SG,in}}+\frac{a^{+}}{n_{\text{H}}\dot{m}\,c_{\text{p,f}}},\quad\text{ and}\\[4.30554pt] T^{\text{HT,in}}(t)&=T^{\text{HT,in}}_{n}&=\tau^{\text{in}}(A_{n})&\text{with}\quad\tau^{\text{in}}(a)&=T^{\text{SG,out}}+\frac{a^{-}}{n_{\text{H}}\dot{m}\,c_{\text{p,f}}},\end{array}$ (8)
@@
with $g^{HF}(\tau^{\text{in}},\tau^{\text{out}})=n_{\text{H}}\dot{m}\,c_{\text{p,f}}(\tau^{\text{out}}-\tau^{\text{in}}-(T^{\text{SG,in}}-T^{\text{SG,out}}))$. Note that $T^{\text{HT,in}}_{n}=T^{\text{SG,out}}$ during charging, $T^{\text{HT,out}}_{n}=T^{\text{SG,in}}$ during discharging, and both equations hold true during idle periods.
@@
From Walden2023 it is known that there is a complex relationship between the HTHX oil outlet temperature $T^{\text{HT,out}}$ and the HTHX oil inlet temperature $T^{\text{HT,in}}$, the HTF mass flow $\dot{m}$, the waste heat air temperature $T^{\text{LT,in}}$ at the LTHX inlet, and the compressor shaft speed $D$, which can be expressed by a surrogate model in terms of a multivariate cubic polynomial $F_{1}$ as $T^{\text{HT,out}}(t)=F_{1}(T^{\text{HT,in}}(t),\dot{m},T^{\text{LT,in}},D(t))$. As we know from (8) that $T^{\text{HT,in}}$ and $T^{\text{HT,out}}$ are stepwise functions, i.e., $T^{\text{HT,out}}(t)=T^{\text{HT,out}}_{n}$ and $T^{\text{HT,in}}(t)=T^{\text{HT,in}}_{n}$ in each interval $t_{n},t_{n+1})$ for $n=0,\ldots,N-1$, and based on the Assumption 2.2, it follows directly that $D$ is also constant between two subsequent discrete time points, so we obtain
@@
More specifically, the compressor shaft speed $D$ can be varied (only at discrete time grid points) to determine the pressure ratio within the HTHP and thus the thermal oil temperature $T^{\text{HT,out}}$ at the HTHX outlet. Details on the definition of the polynomial function $F_{1}$ are provided in Section Appendix C.
@@
Another complex relationship from Walden2023 holds for the electricity consumption $P^{\text{H}}$ of the HTHP. Again, this is expressed by a surrogate model in the form of a multivariate quadratic polynomial $F_{2}$ for $n=0,\ldots,N$ as
@@
The factor $n_{\text{H}}$ in (11) as already mentioned above, denoting the number of HTHPs operating in parallel, takes into account that the surrogate model represents only a single HTHP. The detailed definition of $F_{2}$ can be found in Section Appendix C.
@@
$P^{\text{H}}_{n}=P^{\text{G}}(t)+P^{\text{W}}(t),\quad t\in[t_{n},t_{n+1}),~n=0,\ldots,N-1.$ (12)
For $P^{\text{G}}>0$, electricity is drawn (purchased) from the grid, for $P^{\text{G}}<0$, electricity is fed (sold) into the grid. Note, that the dependence of the WT power $P^{\text{W}}$ on the wind speed $W$ is described by the so-called power curve, $P^{\text{W}}=P_{\text{WT}}(W)$. This is a nonlinear function that grows cubically WANG2019109422 at medium wind speeds until it reaches the rated power. This value is kept constant for higher speeds and is set to zero above the cut-out speed and below a cut-in speed of the turbine. More details are given in Subsection Appendix G.3.
@@
Based on (10) and (11), the HTHP’s electricity consumption $P^{\text{H}}_{n}$ can be determined at each time point $t_{n}$ for a given control $A$, which determines, by (8), the temperatures $T^{\text{HT,out}}_{n}$ and $T^{\text{HT,in}}_{n}$ of the HTF at the outlet and inlet of the HTHX, respectively. Suppose that at time $t_{n}$ the control is set to be $A_{n}=a$, then (8) implies $T^{\text{HT,out}}_{n}=\tau^{\text{out}}(a)$ and $T^{\text{HT,in}}_{n}=\tau^{\text{in}}(a)$, and recall that $\dot{m}$ and $T^{\text{LT,in}}$ are constants. Then, in a first step, the corresponding shaft speed $D_{n}=d$ is determined by solving (10) for the unknown $d$, i.e., $\tau^{\text{out}}(a)=F_{1}(d):=F_{1}(\tau^{\text{in}}(a),\dot{m},T^{\text{LT,in}},d)$. Since $F_{1}$ is a cubic polynomial in $d$, the solution is among the real-valued roots of this polynomial. For the given $F_{1}$ and using the fact that the shaft speed is restricted to values within the interval $d_{\text{min}},d_{\text{max}}]$, defined by technical conditions Walden2023 , there is a unique root $d^{*}=d^{*}(a)$ in this interval. In a second step, the corresponding electricity consumption $P^{\text{H}}_{n}$ is obtained by substituting $d^{*}$ into (11) via
@@
To emphasize the dependence of $P^{\text{H}}$ on the control $A$, we introduce the function $\pi^{\text{H}}$ that maps $A$ to the electricity consumption for $n=0,\ldots,N-1$, i.e.,
$P^{\text{H}}_{n}=\pi^{\text{H}}(A_{n}),\quad\text{with}\quad\pi^{\text{H}}(a)=n_{\text{H}}F_{2}(\tau^{\text{in}}(a),\dot{m},T^{\text{LT,in}},d^{*}(a)).$ (14)
@@ 2.4.3 Thermal Energy Storage Operational Modes
[…]
In charging mode, the excess thermal energy is stored in the sensible TES and increases its medium temperature, while during discharging the cooler HTF with temperature $T^{\text{SG,out}}$ absorbs heat from the TES, and decreases the temperature of the storage medium. For details on the charging and discharging mode, we refer to Section Appendix D. In idle mode, the HTF bypasses the TES completely and obviously implies $T^{\text{HT,out}}_{n}=T^{\text{SG,in}}$ and $T^{\text{HT,in}}_{n}=T^{\text{SG,out}}$, which also follows directly from (86) and (89) given in the Supplementary Material. There, we provide further details, including the dependence of the time-varying charging and discharging factors $l^{\text{C}},l^{\text{D}}$ on the TES temperature $R$ and the chosen control $A_{n}$.
@@ 3.1 State Dynamics
Throughout this paper, all stochastic processes and random variables are supposed to be defined on a filtered probability space $(\Omega,\pazocal{F},\mathbb{F},\mathbb{P})$. In particular, that space carries a two-dimensional Brownian motion $B=(B^{W},B^{S})$, with two independent standard Brownian motions $B^{W},B^{S}$ on $0,t_{\text{E}}]$, which will be used below to drive stochastic differential equations (SDEs) describing the dynamics of $W$ and $S$. The filtration $\mathbb{F}$ is assumed to be generated by $B$, that is, $\mathbb{F}=\mathbb{F}^{B}=(\pazocal{F}^{B}(t))_{t\in[0,t_{\text{E}}]}$ with the $\sigma$-algebras $\pazocal{F}^{B}(t)=\sigma\{B(s),s\leq t\}$, augmented by the $\mathbb{P}$-nullsets, so that $\mathbb{F}$ satisfies the usual assumptions. While, the market price of electricity $S$ can also take negative values, the wind speed $W$ is always non-negative. We therefore replace $W$ with $\log W$ and assume for $t\in[0,t_{\text{E}}]$
$\begin{split}\log W(t)&=\mu_{W}(t)+Y^{W}(t)\quad\text{ and }\quad S(t)=\mu_{S}(t)+Y^{S}(t).\end{split}$ (16)
@@
with mean reversion speeds $\lambda_{W},\lambda_{S}>0$, diffusion coefficients $\sigma_{W},\sigma_{S}>0$ and a constant $c_{W}\geq 0$. Due to the different natures of wind speed and electricity price, we assume that $\lambda_{W}\neq\lambda_{S}$ to simplify our analysis. A positive constant $c_{W}$ leads to a negative correlation of the wind speed $W$ and the price process $S$, see Lemma 3.1 and Proposition 3.2 below. This is often observed in energy markets.
The two SDEs above have analytical solutions that allow us to derive the following closed-form expressions for the joint distribution of the pair of random variables $W$ and $S$ as follows, with the proofs provided in Section Appendix F.
@@
Let $0\leq t_{a}<t_{b}\leq t_{\text{E}}$. Then the solutions of the SDEs (17) on $t_{a},t_{b}]$ to given initial values $Y^{W}(t_{a})=y_{W}$ and $Y^{S}(t_{a})=y_{S}$ with $y_{W},y_{S}\in\mathbb{R}$ are Ornstein-Uhlenbeck processes with
@@
Combining the above result for $t_{a}=t_{n}$ and $t_{b}=t_{n+1}=t_{a}+\Delta t$ with the definitions from (16), and recall the notation $W_{n}=W(t_{n}),S_{n}=S(t_{n})$ for $n=0,\ldots,N$, we obtain the following result for the joint distribution of $(\log W_{n+1},S_{n+1})$ given $(\log W_{n},S_{n})$. This will be useful for the construction of the transition operator and the corresponding transition kernel for the MDP’s state process below in (36).
@@
$\begin{split}m^{W}_{n+1}(w)&=\mu_{W}(t_{n+1})+m_{Y^{W}}(\Delta t,\log w-\mu_{W}(t_{n})),\\ m^{S}_{n+1}(w,s)&=\;\mu_{S}(t_{n+1})\,+m_{Y^{S}}(\Delta t,\log w-\mu_{W}(t_{n}),s-\mu_{S}(t_{n})),\end{split}$ (23)
where $m_{Y^{W}},m_{Y^{S}}$ and $\Sigma_{Y}$ are given in Lemma 3.1.
Note that the wind speed $W$ follows a log-normal distribution because $\log W$ is Gaussian. The above result on the conditional distribution of the pairs $(\log W_{n},S_{n})$ and the fact that the dynamics of the stochastic fluctuations $Y^{W},Y^{S}$ are driven by Brownian motions, i.e., processes with independent increments, can be used to derive a recursion for the discrete-time dynamics of the sequence of these pairs which is driven by a sequence of independent standard normally distributed random vectors.